If you want to be 99% confident of estimating the population mean to within a sampling error of plus or minus 6 and the standard deviation is assumed to be 17 what sample size is​ required?
Answer #1

Solution :

Given that,

Population standard deviation = = 17

Margin of error = E = 6

At 99% confidence level the z is,

= 1 - 99%

= 1 - 0.99 = 0.01

/2 = 0.005

Z/2 = Z0.005 = 2.576

sample size = n = [Z/2* / E] 2

n = [ 2.576 * 17 / 6]2

n = 53.27

Sample size = n = 54